OpenStudy (anonymous):

Find the value of x. Round your answer to the nearest tenth. Answer A. 6.2 cm B. 12.7 cm C. 15.5 cm D. 10.9 cm

OpenStudy (anonymous):



OpenStudy (anonymous):

The angles of a triangle all add to 180. Remember also the law of sines: \[\sin(\alpha)/A = \sin(\beta)/B = \sin(\gamma)/C\]

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

You know one angle is 55. The other is 90. So the other angle added to 55 must equal 90.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

Correct. Your other possible solution is to remember that tangent of an angle is related to the opposite side divided to the adjacent side. \[x = 8.8 / \tan(55)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Oh sorry, no it's not B (12.7cm). I was looking at your answers incorrectly.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Try again, the easiest solution is to use the tangent. Start with what you know. You have an angle, the length opposite of it, and you want the adjacent angle.

OpenStudy (anonymous):

6.2

OpenStudy (anonymous):

That is the correct answer. Yes.

OpenStudy (anonymous):

thanks